Using traditional methods, it takes 11.9 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 9 students and observed that they had a mean of 12.2 hours with a standard deviation of 1.5. A level of significance of 0.05 will be used to determine if the technique performs differently than the traditional method. Assume the population distribution is approximately normal. Find the value of the test statistic. Round your answer to three decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 11.9

Ha : \mu   \neq 11.9

\bar x = 12.2

\mu = 11.9

\sigma = 1.5

n = 9

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (12.2 - 11.9) / 1.5 / \sqrt 9

= 0.6

Test statistic = 0.600